OpenStudy (anonymous):

r^2-9r+14 factor trinomial

OpenStudy (amistre64):

-7(-2) = 14 -7-2 = -9

OpenStudy (amistre64):

r-7; r-2 lol

OpenStudy (anonymous):

ok I kinda see how you got that -7(-2) is 14 then you minus the two is that correct

OpenStudy (amistre64):

there are two different terms to math with the same 2 numbers: the last term is a product of 2 numbers: -7(-2) = 14 the middle term is the sum of the same 2 numbers: -7-2 = -9

OpenStudy (amistre64):

terms to *match.....